### v1

# Chenjie Zhong    Zhipeng Li†,*    Shangzhi Xu Xiaohu Li

Abstract
Variational Mode Decomposition (VMD) plays an important role in many scientific areas, especially for the area of signal processing. Unlike the traditional Fourier paradigm, it makes decomposition of a signal possible without any predefined function basis, which gives unprecedented flexibilities while handling narrow-banded signals of varieties. However, determining the number and central frequencies of intrinsic mode functions are still open questions in that few studies has been proposed to give a complete method that can theoretically guarantee the global convergence during decomposition. In this article, we propose a globally convergent numerical optimization method based on variational convex optimization to automatically determine the number and central frequency of IMFs without any prior knowledge for narrow banded signals, as long as the spectra of each sub-band are identifiable. Our method focuses on finding the support baseline of the spectral function, and further separating the significant frequency band regions above the support baseline. Unlike pioneer works that focus on optimizations on complex field, our method achieves obtaining the number of decomposed IMFs and center frequencies in real field by combining variational calculus, convex optimization, and numerical solutions of differential equations in real field and theoretical analysis shows our algorithm is guaranteed terminate to one of the optimum as close as possible. Experiments also shows that our algorithm converges quickly and can be used in practical engineering to determine the prior information of the number of intrinsic modes and evaluate initial center frequency as prior information to continue the VMD procedure.
keywords:
narrow-banded signals,variational optimization, Intrinsic Mode Function(IMF)
{MSCcodes}
65K10, 65L10, 68U01, 46N10, 34B05
1 Introduction
Variational Mode Decomposition (VMD) [5] decomposes original signals into multiple Intrinsic Mode Functions (IMF) with limited bandwidth with the help of variational optimization. The core of VMD is to minimize the sum of the bandwidths with respect to the estimated IMFs as well as to ensure the sum of all the IMFs is equal to the original signal. The paradigm of VMD is a significant progress in the field of signal processing in that it not only inherits the core of EMD that breaks the traditional Fourier-transform alike framework for signal decomposition where a signal must be decomposed in terms of fixed signal basis like sine and cosine functions, but is also considered as the first trial to solve the signal decomposition problem with the help of variational optimization. VMD plays an important role in modern signal decomposition after proposition. In the field of mechanics, [14] uses VMD to decompose the vibration signals of planetary gearboxes under variable-speed conditions, in biomedical field, VMD can be used to diagnose the seizure of epileptic by retrieving the feature of decomposed Electroencephalography (EEG) signals [21] and can also help to denoise the Electrocardiogram (ECG) signals [9]. In the field of Earth Physics signals, it not only helps to increase the Signal-to-Noise Ratio(SNR) by combining with adaptive non-local mean filtering [7], but also helps the prediction of monthly runoff by decomposing the signal into trend, periodic and random components as the input feature of downside neural networks [16]. Besides, VMD also provides the potential for the channel identification in communications [1] and helps to the visualization of the protein structure [15]. We show the detail of VMD in Supplementary Materials Section 1. Although VMD has provided flexibility in decomposing signals, there are factors that hinder its further applications in that the number of intrinsic mode function must be determined manually as a piror [19, 10, 18, 17] in coordination with penalty factor without a common principle. Besides, VMD procedure works in complex field and to our knowledge there is no rigorous analysis on the convergence[5]. In order to alleviate the problem, the subsequent research either needs the knowledge to the range of the number of modes [11], or retrieves the IMFs recursively [12], or highly dependent on a predefined window-size to retrieve the peak and valley of spectra [6]. To our best knowledge, although there are some research focusing on wide-band extension of VMD [22], yet even to narrow-band signals, there are almost no complete research that can simultaneously be rigorous in theoretical analysis, while obtaining estimation of the number of modes and center frequency without any prior knowledge to the signal, and achieving non-recursive iterations.
In this article, we propose a variation-optimization based, globally convergent method to automatically determine the IMFs numbers and the corresponding center frequencies for narrow-banded signals. In short, we model the modes retrieving problem as adaptively extracting the supporting baseline of the signal spectrum amplitude function. By subtracting the extracted baseline from the signal spectrum amplitude function, we retrieve the significant frequency band components of the signal and extract the various modes of the narrow-band signal. The extracted supporting baseline is acquired by adversarially maximizing the integral of the baseline while minimizing the integral of its curvature. The supporting baseline is computed by iteratively solving a variational problem that is degenerated into the solution of a normal ordinary differential equation, and under proper condition and weight parameters the method always converges to the optimal curve that can be used to determine the center frequencies and the number of IMFs after proper estimation of the gap residual at the bottom. Once the number of IMFs and initialized frequencies are determined by our method, these parameters can be used to start the following VMD procedure. Experiments show that our method can quickly get accurate estimation about the number of IMFs and the corresponding initial center frequencies as prior to the following VMD procedure.
Our contribution can be summarized as 3 folds. First, as far as we know, in determining the number and central frequencies of IMFs, we are the first one to not only give a feasible algorithm that can decompose a narrow-banded signal automatically but also prove the global convergence in a rigorous mathematical view. Second, our proof generalizes the classical dual ascent method into functionals, which expands the scope of the application for the classical algorithm. Third, experiments show that our algorithm can filter out all the potential modes without any prior setting or assumptions to the number of modes, and also show good performance in both precision and speed even when the center frequencies of different modes are very close.
This article will be organized as follows: In the following section we propose the main framework of our algorithm and give the proof of the convergence. In section 3 we introduce the implementation detail about the whole workflow. In Section 4 we design a series of experiments to evaluate the performance of our algorithm. In the final section we show our conclusion about the main contribution of the paper.
2 Methodology
In this section we introduce our method for estimating the number of IMF modes as well as center frequencies of each band. We first introduce the background of the problem and give a brief introduction about Variational Mode Decomposition, and then we propose the concept of Supporting Baseline of a Function, finally we describe our method for finding the baseline and the whole workflow.
2.1 Supporting Baseline of a Function
Without loss of generality we first introduce the Supporting Baseline of a Function. For any function $s(x)>0$, the supporting baseline is a curve that tightly fits the lower bound of the function and reflects the overall trend of the lower bound. Fig. 1 gives an intuitive illustration about the conception of supporting baseline. As illustrated in Fig. 1, intuitively the supporting baseline should be tightly close to the lower bound of the function, approaching the infimum of functions as close as possible as well as maintain its own smoothness in order to capture the global trend of the lower bound rather than too much details, as capturing too much details will in turn reduce the robustness for separating the modes since there might be some noise added to the signal in either time or frequency domain.
Figure 1: Illustration to the concept of supporting baseline: In (a)-(c) the solid line is $y=10\cos\pi x-10\cos 5\pi x+x^{2}$. The dashed line in (a) is a proper supporting baseline since it actually fits the lower bound of the solid line as well as ignoring high frequent details of the solid line. The dashed line in (b) has a gap to the bottom of the solid line so it is not suitable to be a good supporting baseline. In contrast, the dashed line in (c) fits too tightly to the lower bound of the solid line, capturing too much detail of the solid line so that it can be reckoned as the lower envelope rather than a proper supporting baseline.
2.2 Variational Methods for Finding the Supporting Baseline
To solve the problem for finding the supporting baseline, we use optimization method mainly based on variation and convex optimization [8, 4]. Let $f(x)\geq 0$ be a known function, the problem of finding the supporting baseline curve of $f(x)$ is equivalent to finding a function $g(x)$ that tightly approaching the infimum of $f(x)$ while maintaining the global trend of its lower bound. This can be viewed as the following optimization problem:
[…]
Converting (1) into standard form, we get:
[…]
This is an optimization problem with inequality constraints. To solve the problem, one can construct an Lagrange Slack Function by adding functional-styled weights to each factor and constraints
$L(g,\lambda,\mu,\alpha,\beta)$ $=\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x+\int_{\Omega}\beta(x)[-g(x)]\text{d}x$
(3) $+\int_{\Omega}\lambda(x)[g(x)-f(x)]\text{d}x+\int_{\Omega}\mu(x)[-g(x)]\text{d}x$
where
[…]
and $\alpha(x),\beta(x)>0$. The solution of the problem can be found by solving a variation problem, namely, if we set
(6) $F=\alpha(x)g^{\prime\prime 2}(x)-\beta(x)g(x)+\lambda(x)[g(x)-f(x)]-\mu(x)g(x)$
then the above Lagrange Slack Function is like the form $J[g(x)]=\int_{\Omega}F(x,g,g^{\prime\prime})\text{d}x$ so that if we assume the variation is zero at both ends of the curve, one can easily know the optimal solution of the above function happen to satisfy the simplified form of Euler-Poisson equation [8] if we let the variation vanish, so it can be written as
(7) $F_{g}+\cfrac{\text{d}^{2}F_{g^{\prime\prime}}}{\text{d}x^{2}}=0$
where
(8) $F_{g}=-\beta(x)+\lambda(x)-\mu(x)$
(9) $F_{g^{\prime\prime}}=2\alpha(x)g^{\prime\prime}(x)$
It is easily to get that
(10) $\cfrac{\text{d}^{2}F_{g^{\prime\prime}}}{\text{d}x^{2}}=2\alpha^{\prime\prime}(x)g^{\prime\prime}(x)+4\alpha^{\prime}(x)g^{(3)}(x)+2\alpha(x)g^{(4)}(x)$
By substituting (8) and (10) into (7), we can get the following ordinary differential equation
(11) $2\alpha^{\prime\prime}(x)g^{\prime\prime}(x)+4\alpha^{\prime}(x)g^{(3)}(x)+2\alpha(x)g^{(4)}(x)=\beta(x)-\lambda(x)+\mu(x)$
This is a 4th-ordered Ordinary Differential Equation (ODE), so that normally four constraints are needed to determine a definite solution. Recall that we assume the variation is zero at both ends of the curve, so that we can typically set four constraints at both ends of the curve with respect to the function value and its first-order derivative as follows:
(12) $g(0)=f(0),g(n)=f(n),g^{\prime}(0)=f^{\prime}(0),g^{\prime}(n)=f^{\prime}(n)$
where 0 stands for the left end and $n$ the right end of the curve. To solve the boundary value problem of ODE, we use Finite Difference Method [20]. Note that for a differential function $g(x)$, it’s derivative of first to fourth order can be expressed approximately as the linear combination of $g(x)$ itself and we can take similar operation to the function $\alpha(x)$ (See Supplementary Materials 2). Since now different order of derivative with respect to $\alpha(x)$ and $g(x)$ can be expressed as the linear combination of themselves, one can easily transform (11) into the matrix-form as follows:
(13) $(2\boldsymbol{A^{(2)}\alpha})\odot(\boldsymbol{G^{(2)}g})+(4\boldsymbol{A^{(1)}\alpha})\odot(\boldsymbol{G^{(3)}g})+(2\boldsymbol{\alpha})\odot(\boldsymbol{G^{(4)}g})=\boldsymbol{\beta-\lambda+\mu}$
where $\boldsymbol{A}^{(n)}$ and $\boldsymbol{G}^{(n)}$ is the conversion matrices (See Supplementary Materials 2) that transform the nth derivatives of $\alpha(x)$ and $g(x)$ into linear expression of the corresponding function, together with the boundary condition to form full-rank matrices, and $\boldsymbol{\alpha}$, $\boldsymbol{g}$, $\boldsymbol{\beta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\mu}$ is obtained by respectively discretizing the corresponding functions, $\odot$ stands for component-wise multiplication between matrices(See Supplementary Materials 3 for details). The conversion matrices $\boldsymbol{A}^{(n)}$ and $\boldsymbol{G}^{(n)}$ share the same structure, in which the last four lines depicts the boundary condition and the other the ODE equation, and $\boldsymbol{g}$, $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\mu}$ are column vectors in which each element is a uniformly discrete sampled value on the function $g(x)$, $\alpha(x)$, $\beta(x)$, $\lambda(x)$, $\mu(x)$, respectively.
2.3 Proof of the Convergence
Algorithm 1 The Projection Gradient Descent Method for Evaluating Support Baseline
1: Input $\{f(x)\}$
2: Initialize $\alpha(x),\beta(x),\lambda(x),\mu(x),g(x)$
3: Pre-process by extending $f(x)$ at both ends.
4:repeat
5:  $n\leftarrow n+1$
6:  Update $g(x)$:
7:   Solve the equation:
8:   $(2\boldsymbol{A^{(2)}\alpha})\odot(\boldsymbol{G^{(2)}g})+(4\boldsymbol{A^{(1)}\alpha})\odot(\boldsymbol{G^{(3)}g})+(2\boldsymbol{\alpha})\odot(\boldsymbol{G^{(4)}g})=\boldsymbol{\beta-\lambda+\mu}$
9:  Dual ascent:
10:   $\lambda^{n+1}\leftarrow\max\{0,\lambda^{n}+\theta(g-f)\}$
11:   $\mu^{n+1}\leftarrow\max\{0,\mu^{n}+\gamma(-g)\}$
12:until convergence: $\|g^{n}-g^{n-1}\|_{2}^{2}/\|g^{n}\|_{2}^{2}<\epsilon$
In this section we show the global convergence of the Algorithm 1. In our algorithm we have in fact solved a functional Extreme Value Problem with respect to the spectrum to the original signal. Despite the fact that the Fourier transform spectrum of a general signal may not necessarily be smooth, yet on one side, as in practical digital signal processing, the signal are discretized and can be smoothly interpolated, indicating the function processed by our algorithm are practically smooth which enable the reasonable use of commonly used assumptions that are advantageous for proving algorithm convergence, and on the other hand, subsequent numerical experiments have also shown that such assumptions not only do not affect the range of signals that can be processed, but are also very helpful for proving convergence. Therefore, in our proof, we assume that the spectrum is smooth so that we can extend the Lagrangian dual gradient ascent to functionals. Our proof combines functional, variational optimization and ODE theories in the framework of gradient projection, and can be divided into 3 steps. First we prove that we are dealing with a convex functional optimization problem, where the feasible domain is also a convex set. Next we show that under the condition of our algorithm, the Lagrangian dual gap is zero [4, 2] so that it is possible to find the optimum by dual ascent method with some modification from vectors to functional. After that, the convergence of iteration in dual space is shown and it is also shown that the optimal function can also be approached during the iterations by utilizing the dependence of differential equation solutions on coefficients. For the sake of limited spaces, we move the proof of convexity and duality into Supplementary Materials 4 and only give the proof to the convergence of iteration in Algorithm 1 in this part.
We first restate the iteration in Algorithm 1. Notice that during the iteration, we have
(14) $\left\{\begin{aligned} g_{k+1}(x)&=\underset{g(x)}{\arg\min}\mathcal{L}(g(x),\lambda_{k}(x),\mu_{k}(x))\\ \lambda_{k+1}(x)&=\max[0,\lambda_{k}(x)+\theta(g_{k+1}(x)-f(x))]\\ \mu_{k+1}(x)&=\max[0,\mu_{k}(x)+\gamma(-g_{k+1}(x))]\end{aligned}\right$.
to show the projection property for the $\max$ operation we first prove the following lemma:
Lemma 2.1.
Let $m(x),x\in\mathcal{D}$ be any continuous function that has both positive and negative values, Let $\Pi$ be an operator that makes all the negative values for $m(x)$ zero and non-negative unchanged, or $\Pi m(x)=\max(0,m(x))$, let $M(x),x\in\mathcal{D}$ be any non-negative function, then
(15) $\int_{\mathcal{D}}\bigl[m(x)-\Pi m(x)\bigr]\bigl[M(x)-\Pi m(x)\bigr]dx\leq 0$
Proof 2.2.
By expanding the equation (15) we should have the following:
(16) $\begin{split}&\ \ \ \ \int_{\mathcal{D}}\bigl[m(x)-\Pi m(x)\bigr]\bigl[M(x)-\Pi m(x)\bigr]dx\\ &=\int_{\mathcal{D}}m(x)M(x)dx-\int_{\mathcal{D}}\Pi m(x)M(x)dx-\int_{\mathcal{D}}m(x)\Pi m(x)dx+\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx\\ &=\int_{\mathcal{D}}m(x)M(x)dx-\int_{\mathcal{D}}\Pi m(x)M(x)dx-\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx+\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx\\ &\leq\int_{\mathcal{D}}m(x)M(x)dx-\int_{\mathcal{D}}m(x)M(x)dx-\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx+\int_{\mathcal{D}}\bigl[\Pi m(x)\bigr]^{2}dx=0\end{split}$
Moreover, if $m(x)$ is in feasible set so that $m(x)=\Pi m(x)$, then (16) will become zero.
By using Lemma 15, we can prove that in iteration (14) the multiplier $\lambda_{k}(x)$, $\mu_{k}(x)$ always terminate at one of the optimum. To say in detail, we should have the following theorem:
Theorem 2.3.
Let $\mathcal{G}(\lambda(x),\mu(x))=\underset{g(x)}{\arg\min}\mathcal{L}(g(x),\lambda(x),\mu(x))$ and $\lambda_{k}(x)$, $\mu_{k}(x)$ be equicontinuous for all $k=1,2,3,\cdots$, then if $\mathcal{G}$ satisfies
[…]
for any $\lambda_{1}(x)$, $\lambda_{2}(x)$, $\mu_{1}(x)$, $\mu_{2}(x)$, where $D$ stands for the variation, and the step size $\eta<\frac{1}{L}$, then as $k\rightarrow+\infty$, the iteration series of $(\lambda_{k}(x),\mu_{k}(x))$ must terminate at the point as close as possible to one of the optimum $(\lambda^{*}(x),\mu^{*}(x))$ for $\mathcal{G}$.
Proof 2.4.
For conciseness, simplifying the integral by introducing matrix-form expression, where
(21) $\left[\begin{array}[]{cc}a(x)&b(x)\end{array}\right]\left[\begin{array}[]{cc}c(x)\\ d(x)\end{array}\right]=\int_{\mathcal{D}}a(x)c(x)\text{d}x+\int_{\mathcal{D}}b(x)d(x)\text{d}x$
we have
(25) $\left[\begin{array}[]{cc}\tilde{\lambda}_{k+1}(x)-\lambda_{k+1}(x)&\tilde{\mu}_{k+1}(x)-\mu_{k+1}(x)\end{array}\right]\left[\begin{array}[]{cc}\lambda_{k}(x)-\lambda_{k+1}(x)\\ \mu_{k}(x)-\mu_{k+1}(x)\end{array}\right]\leq 0$
[…]
On the other side, for any functional we have
$\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))=\mathcal{G}(\lambda_{k}(x),\mu_{k}(x))$
$+DG_{(\lambda_{k}(x),\mu_{k}(x))}(\lambda(x),\mu(x))^{\top}\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}$
$+\frac{1}{2!}\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)&\mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}$
(31) $\quad D^{2}G_{(\lambda(x),\mu(x))}(\varepsilon_{k}(x),\zeta_{k}(x))\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}$
where $(\varepsilon_{k}(x),\zeta_{k}(x))$ locates on general line segment from $(\lambda_{k}(x),\mu_{k}(x))$ to
$(\lambda_{k+1}(x),\mu_{k+1}(x))$. If $\mathcal{G}$ is second-order differentiable and satisfies (17) so that $D^{2}\mathcal{G}$ is bounded with [L,L][-L,L]. Due to the concavity of $\mathcal{G}$ and the left side of the bound, and combine (30) we get
[…]
if $\eta=\max\{\theta,\gamma\}=\cfrac{1}{L}<\cfrac{2}{L}$, then $\mathcal{G}$ is a monotonously increasing functional series with upper bound due to its concavity, so that $\mathcal{G}$ converges and the difference between adjacent term of both $\mathcal{G}$ and $(\lambda_{k}(x),\mu_{k}(x))$ is approaching zero.
Now we prove $\{g_{k}(x)\}$ is guaranteed to approach one of the optimum when the algorithm terminates even if $\mathcal{G}$ is not strongly concave. Due to the concavity the iteration on $\mathcal{G}$ should satisfy [2]
$l\left(\begin{bmatrix}\lambda_{k+1}(x)\\ \mu_{k+1}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)-\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda_{k+1}(x)-\lambda_{k}(x)\\ \mu_{k+1}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
$\geq l\left(\begin{bmatrix}\lambda^{*}(x)\\ \mu^{*}(x)\end{bmatrix};\begin{bmatrix}\lambda_{k}(x)\\ \mu_{k}(x)\end{bmatrix}\right)-\cfrac{1}{2\eta}\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
(35) $+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}$
since $\mathcal{G}$ is convex. Note that the spectrum of $D^{2}\mathcal{G}$ is larger than L-L and $\eta=\cfrac{1}{L}$ to replace the left side of (35), and use the concavity of $\mathcal{G}$, we have
$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq 2\eta(\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))$
(36) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)))+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
Sum up (36) for each $k$ and eliminate the repeated terms that occur at the both ends of the inequality, we have
$\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}\leq 2\eta\sum\limits_{i=0}^{k}(\mathcal{G}(\lambda_{i+1}(x),\mu_{i+1}(x))$
(37) $-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)))+\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{0}(x)\\ \mu^{*}(x)-\mu_{0}(x)\end{bmatrix}\right\|_{2}^{2}$
As the left side is non-negative, and recall in (34) we know that $\{\mathcal{G}_{k}\}$ is non-decreasing and $\mathcal{G}^{*}$ is the supreme of $\mathcal{G}$, hence $\mathcal{G}(\lambda_{k+1}(x),\mu_{k+1}(x))-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))$ is the largest among all the terms involved in the summation, by applying reasonable inequality scaling we must have
(38) $0\geq\mathcal{G}(\lambda_{k+1}(x)$ $,\mu_{k+1}(x))-\mathcal{G}(\lambda^{*}(x),\mu^{*}(x))\geq-\cfrac{1}{2\eta(k+1)}\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{0}(x)\\ \mu^{*}(x)-\mu_{0}(x)\end{bmatrix}\right\|_{2}^{2}$
which shows the iteration is approaching the optimum. From (36) and by utilizing the monotonicity of the functional $\mathcal{G}$ during iteration we have
(39) $\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k+1}(x)\\ \mu^{*}(x)-\mu_{k+1}(x)\end{bmatrix}\right\|_{2}^{2}<\left\|\begin{bmatrix}\lambda^{*}(x)-\lambda_{k}(x)\\ \mu^{*}(x)-\mu_{k}(x)\end{bmatrix}\right\|_{2}^{2}$
which shows the distance between the iterating points and the optimum is always decreasing, since the distance is always non-negative so that it has lower bound, $(\lambda_{k}(x),\mu_{k}(x))-(\lambda^{*}(x),\mu^{*}(x))$ is a convergent sequence so that $(\lambda_{k}(x),\mu_{k}(x))$ is bounded. According to Arzelà-Ascoli Theorem, there exists some sub-sequences $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$ that converges, so that we cannot assert that the sequence is necessarily convergent in all cases and need careful discussion. One one hand, if there is only one optimum in $\mathcal{G}$, then it is trivial to show that all these sub-sequences converges to the unique optimum by the constraints in (38). If $\mathcal{G}$ has more than one optimum, then suppose there exists more than one sub-sequence, for example, $(\lambda_{k_{i}}(x),\mu_{k_{i}}(x))$, $(\lambda_{k_{j}}(x),\mu_{k_{j}}(x))$ that converges to $(\lambda^{*}_{i}(x),\mu^{*}_{i}(x))$, $(\lambda^{*}_{j}(x),\mu^{*}_{j}(x))$ respectively, then due to (34) we show that although the sequence might have converged subsequence that approaches to different optimum, the adjacent term tends to zero so that the adjacent primal solution in iteration also tends to zero. On the other hand, as numerical algorithms will introduce error inevitably, if after some iteration $(\lambda_{k}(x),\mu_{k}(x))$ is exactly the unique optimum so that the right hand of (39) is zero, then from monotonicity of iteration we know $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$. Even if $\mathcal{G}$ has more than one optimum and after some iteration $(\lambda_{k}(x),\mu_{k}(x))$ approaches one of the optima $(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))$, since there is another optimum $(\lambda^{*}_{2}(x),\mu^{*}_{2}(x))$, then we have
(40) $\mathcal{G}(m\lambda^{*}_{1}(x)$ $+(1-m)\lambda^{*}_{2}(x),m\mu^{*}_{1}(x)+(1-m)\mu^{*}_{2}(x))$
$=\mathcal{G}(\lambda^{*}_{1}(x),\mu^{*}_{1}(x))=\mathcal{G}(\lambda^{*}_{2}(x),\mu^{*}_{2}(x)),m\in[0,1]$
since the direction of the line connecting the median point and the initial point in the first-order Taylor expansion formula is consistent with $\left[\begin{array}[]{c}\lambda^{*}_{2}(x)-\lambda^{*}_{1}(x)\\ \mu^{*}_{2}(x)-\mu^{*}_{1}(x)\end{array}\right]$, so we have
(41) $D\mathcal{G}_{(\lambda_{k}(x),\mu_{k}(x))}(\lambda(x),\mu(x))\left[\begin{array}[]{c}\lambda^{*}_{2}(x)-\lambda^{*}_{1}(x)\\ \mu^{*}_{2}(x)-\mu^{*}_{1}(x)\end{array}\right]=0$
thus the variation direction is always orthogonal to any lines connected between different optima, as a consequence, when projected again during the next iteration, $(\lambda_{k+1}(x),\mu_{k+1}(x))$ must overlap with $(\lambda_{k}(x),\mu_{k}(x))$, so we have $(\lambda_{k+1}(x),\mu_{k+1}(x))=(\lambda_{k}(x),\mu_{k}(x))$. In all cases we show $(\lambda_{k}(x),\mu_{k}(x))$ is always convergent to at least one of the optimum and stay still as long as it arrives the optimum during the iteration even if there are more than one optimum in $\mathcal{G}$.
Now we turn to prove the convergence of series $g_{k}(x)$. As each $g_{k}(x)$ is obtained by solving an ODE function as the optimum with respect to a variational problem, we should first prove the existence and uniqueness to the solution of $g$ for each iteration.
Theorem 2.5.
The solution of (11) with $g(0)=0,g(n)=0,g^{\prime}(0)=0,g^{\prime}(n)=0$ is $g(x)\equiv 0$.
Proof 2.6.
Note that the differential equation of the boundary value problem we are solving has only zero solutions under the corresponding homogeneous differential equation and homogeneous boundary conditions, in order to prove the existence and uniqueness of our solution in each iteration [3]. Recall that when dealing with the homogeneous condition, we exactly have
(42) $g(x)=\int_{a}^{x}\left(\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t\right)\text{d}t+C_{3}x+C_{4}$
where
(43) $g(a)=g(b)=g^{\prime}(a)=g^{\prime}(b)=0$
It is obvious that $C_{3}$ and $C_{4}$ is zero by $g(a)$ and $g^{\prime}(a)$, and $g^{\prime}(b)=0$ is equivalent to
(44) $\int_{a}^{b}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t=0$
If $\alpha(t)>0$ and $\alpha(t)\in C[a,b]$, since $C_{1}t+C_{2}$ is a linear function so that if the zero point is not in $[a,b]$ then it is not possible to make (44) hold. If there exists $t\in[a,b]$ that makes $C_{1}t+C_{2}$ zero, then the upper limit integral function $\Phi(x)=\displaystyle\int_{a}^{x}\cfrac{C_{1}t+C_{2}}{\alpha(t)}\text{d}t$ has no zero points in $(a,b)$, thus makes $g(b)=\displaystyle\int_{a}^{b}\Phi(t)\text{d}t$ non-zero, which is paradoxical to our assumption if not both $C_{1}$ and $C_{2}$ is zero. As a consequence, during each iteration, there always exist a unique solution to $g(x)$, so that we can simply discuss the convergence based on the continuous-dependency between the solution and the coefficients.
Corollary 2.7.
For each iteration of (13), there always exists a unique solution.
Now we begin to prove the convergence of $g_{k}(x)$. For this property during the iteration in Algorithm 1, we have the following theorem:
Theorem 2.8.
Let $f(x)$ be first-ordered Lipschitz-continuous function that satisfies $|f^{\prime}(x)-f^{\prime}(y)|\leq L|x-y|$ for any $x,y$, and suppose $g_{0}(x)$ satisfies $0\leq g_{0}(x)\leq f(x)$. If $\{g_{k}(x)\}$ satisfies
[…]
Then if the step $\eta$ is sufficiently small so that $\eta<\frac{1}{L}$ for each iteration, $g_{k}(x)$ must terminate at one of the optimum.
Proof 2.9.
Based on lemma 15, theorem 2.3 and theorem 2.5 we have shown that our original problem has unique solution during each iteration. As (11) is a linear ODE, its solution is continuously dependent on its coefficients, so we have $\lim_{k\rightarrow+\infty}g_{k}(x)=g^{*}(x)$ when $\{(\lambda_{k}(x),\mu_{k}(x))\}$ converges or falls early in one of the optima, and
[…]
$\{(\lambda_{k}(x),\mu_{k}(x))\}$ has sub-sequences approaching different optimum respectively. In all cases we can terminate algorithm and guarantee the $\{g_{k}(x)\}$ terminates to one of the optimum as close as possible.
3 Implementation
3.1 Algorithm
In this part we state the whole implementation of our method. Our implementation can be mainly divided into three parts. Given a signal to be decomposed, we first use variational method in terms of iteratively constructing such a function sequence that each term of it is the optimal solution obtained by vanishing the variation of the Lagrangian Slack Function under the current evolutionary state with respect to inequality constraint conditions and then updating the weight of constraint using dual ascent method. Thirdly, we apply Kernel Density Estimation (KDE) to the amplitude values of each point in the spectrum to eliminate residual gaps at the bottom, which is based on the observation that the residual gap at the bottom will results in a clustering behaviour around the corresponding value that the residual gap share. The proper threshold to eliminate the gap for further evaluating the number of IMFs should locate where the value of the KDE function achieves the maximum. In short, our whole workflow can be summarized as in Algorithm 2.
Algorithm 2 The Whole Workflow
1: Input $\{f(x)\}$
2: Initialize $\alpha(x),\beta(x),\lambda(x),\mu(x),g(x)$
3: Pre-process by extending $f(x)$ at both ends.
4:repeat
5:  Run the kernel process of Algorithm 1
6:until convergence: $\|g^{n}-g^{n-1}\|_{2}^{2}/\|g^{n}\|_{2}^{2}<\epsilon$
7: Estimate KDE for the computed $g(x)$.
8: Compute best threshold $t_{best}=\arg\max\text{KDE}_{g(x)}(t)$.
9: Subtract noise-gap threshold from $g(x)$ by $g(x)-t_{\text{best}}$.
10: Get $k$, $\{\omega_{k}\}$ from non-zero intervals.
11: Begin VMD procedure [5] with $k$ and $\{\omega_{k}\}$
12: Output $\{u_{k}\}$
3.2 Pre-processing
In order to solve the equation effectively, before we start the iteration, we first normalize the signal into $[0,1]$, in both domain and range. Since in each iteration, our optimal baseline should satisfy the boundary condition of the ODE, so that when dealing with high-pass or low-pass signals where spectrum has sharp slope at the ends, the to-be-evaluated baseline will also exhibits the sharp slope, which will in turn have bad effect on the convergence at the ends. To overcome the boundary effect, we extend our original function by making a gentle decay at both ends of the spectrum that not only completes the peaks at both ends but also alleviate the sharp slope effect. We will discuss this in detail later in our experiments.
3.3 Post-processing
Despite the fact that the iteration procedure to find the supporting baseline converges in reasonable time in most cases when applied with proper extension to the original spectrum of the signal, the shape of the spectrum do have great impact on the convergence rate. For example, since our optimization strategy discourage the curvature (via second-order derivative) of the curve, the optimal baseline is prone to converge more faster at points where original spectrum is flat. On the other side, although $g(x)$ will eventually adhere closely to the lower bound of $f(x)$, it is difficult to guarantee that the error between $f(x)$ and $g(x)$ calculated after stopping the iteration is consistent at every point so that f(x)g(x)f(x)-g(x) will have some non-zero points where it should be zero, which makes further retrieving modes of difficulty. Since $g(x)$ converges to the bottom of $f(x)$, it is reasonable to assume that the gap is typically small so that f(x)g(x)f(x)-g(x) is almost zero except for those frequency points corresponds with effective signal components. In order to eliminate the gap, we use Kernel Density Estimation(KDE) method to estimate the best threshold, so that the gap of f(x)g(x)f(x)-g(x) will be exposed where the value of the KDE function achieves the maximum. Fig. 2 illustrates such a phenomenon.
Figure 2: Illustration of the non-consistence of the error between the converged optimal supporting baseline and the original function. Fig. 2a shows the converged $g(x)$ and Fig. 2b shows f(x)g(x)f(x)-g(x). Note that since $g(x)$ does have gap to the lower bound of $f(x)$ so that the bottom of f(x)g(x)f(x)-g(x) is almost not on a horizontal line everywhere, which indicates there are some non-zero gaps. These non-zero gaps will have a significant impact on the stability of the subsequent modal number solution. 2c shows the estimate kernel density pertaining the gap at the bottom. 2d shows the estimated f(x)g(x)f(x)-g(x) after removal of the gap at the bottom.
4 Numerical Results
In this section we first show our experiments to our method. In order to prove the effectiveness of our method, we choose some standard function that were repeatedly introduced as benchmarks in some pioneer works [5, 13, 12], yet pay more attention on investigating the evolution of our algorithm during the course of iteration, to show the convergence and robustness of our method. Since our algorithm can be reckoned as an algorithm for finding the number of intrinsic modes and initial center frequencies that the proceeding VMD module needs, our experimental focus is on exploring the behaviour of the algorithm towards various signals in the process of finding the baseline and talking about the KDE process for eliminating the gap, rather than elaborating the performance brought by the subsequent VMD decomposition in the final decomposition results. Then we turn to discuss why we should and how we eliminate the boundary effect by extending the spectrum obtained from the FFT of original signals.
4.1 Experiments on Finding the Supporting Baseline
In this part we show our experiments on finding the supporting baseline for some benchmark signals that were frequently used to test the performance of algorithms in several articles [5, 12]. As default, all signals are obtained by linearly normalizing the domain of the original function to between 0 and 1, and the range of which is normalized into 0 and 1 before iterations to keep the stability. We uniformly add 200 points to the domain and only preserve half of the transformed spectrum as our initial spectrum to be iterated according to the properties of Fourier Transform with respect to real signals. In this way, the Nyquist frequency is 100Hz, which means the highest signal angular frequency that can be effectively represented is $200\pi$, which corresponds to 1 in our normalized FFT results. Our convergence criterion is $\cfrac{\|g_{n}-g_{n-1}\|}{\|g_{n}\|}<10^{-4}$.
We choose five typical signals including single mode signal, multi-mode signal, piecewise continuous signal, signal that are not necessarily narrow-banded, composite signal in which the center frequency of its components are very close. Due to the space limitations, we only illustrate the final convergent state of the iteration for the evaluation of the support baseline, as well as the convergent curve with respect to the error, and indicate the number of iteration steps in the caption. The intermediate iteration process status and the decomposed signal is given in detail in the supplementary materials 5.
Experiment 1: The first example is a single mode signal, on the other words, the signal is narrow-banded and only contains one center frequency. Here we choose $y(t)=100\sin 20\pi t$. As we see the normalized center frequency is $0.1$.
Figure 3: The initial, final convergent state and convergence curve in logarithmic coordinate system with respect to the error for finding the optimum supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$. The iteration lasts for 5161 steps and costs 1.88 seconds.
Experiment 2: The second example is a multi-mode signal, $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$, which has two modes, exactly. The normalized center frequency is 0.05 and 0.1, which is relatively very close to each other. However, our algorithm can still survive where center frequencies of different modes are very close, since from 6000th iteration and on, the support baseline forms a small bulge at the close frequency points, which just separates the two frequency points.
Figure 4: The final convergent state in logarithmic coordinate system for finding the supporting baseline of FFT with respect to FFT of $y(t)=10\cos 10\pi t+20\sin 20\pi t$. The iteration lasts for 5258 steps and costs 1.89 seconds.
Experiment 3: The next example is a little complicated, it has four modes totally, we choose
(45) $\footnotesize y(t)=\left\{\begin{aligned} 6t^{2}+\cos(10\pi t+10\pi t^{2})+\cos(60\pi t),&\qquad t\in[0,0.5]\\ 6t^{2}+\cos(10\pi t+10\pi t^{2})+\cos(80\pi t-10\pi),&\qquad t\in(0.5,1]\end{aligned}\right$.
as our signal. It has a low frequency component, $6t^{2}$, two pure harmonic components, $\cos 60\pi t,\cos(80\pi t-10\pi)$, in different intervals that does not intersect with each other, respectively, and one narrow banded component, $\cos(10\pi t+10\pi t^{2})$, in the whole interval. To our surprise, the iteration converges even more quickly than those signals that has fewer modes. In our cases, it takes about 15000 iterations to achieve the convergence.
Figure 5: The final convergent state in logarithmic coordinate system for finding the supporting baseline of FFT with respect to time series (45). Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out. The iteration lasts for 3026 steps and costs 1.07 seconds.
Experiment 4: Now we show some examples on signals that conflict with the narrow-band assumption. We take $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. As we see that, the spectrum of the signal exhibits two main peaks, however the signal is not a completely narrow-band signal since it exhibits comb-shaped spectrum with respect to harmonic components, although the energy of those harmonic components are relatively very small(Fig. 6). Our algorithm extracts three components of the signal, despite that in the mainstream view there should be only two IMFs. However, since it is not a narrow-band signal, the difficulty of decomposition objectively exists since one can not even provide a specific standard to determine whether these small harmonic components should be considered as an intrinsic mode. We also observe that, even in [5, 12], when carefully zoomed in, there does exist small oscillations in the first decomposed component which indicates the non-purity of it. Despite the fact that our results has three components, every separate component exhibits better purity than the pioneer works. On the other side, from the perspective of reconstruction error, it is shown in this example that, results generated by algorithm in [5] with both 2 and 3 modes has reconstruction error as twice as large as ours (0.1 ours vs 0.2 theirs, see S5.13a), which also indicates that our algorithm has better performance in this case.
Figure 6: The final convergent state in logarithmic coordinate system for finding the baseline of function $\frac{1}{1.2+\cos 2\pi t}+\frac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. The iteration lasts for 3463 steps and costs 1.25 seconds.
Experiment 5: Next we show even when the central frequencies are very close, our algorithm can also work well. We take signals as $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$. The signal has one low-frequency component and 10 alternating current components with different amplitudes, of which the minimum difference between frequency gaps is only 0.04 after normalization and our algorithm can still survive in figuring out all the components.
Figure 7: The final convergent state in logarithmic coordinate system for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The iteration lasts for 1768 steps and costs 0.64 seconds.
4.2 Experiments on Extension of Original Spectrum at Both Ends
Although supporting baseline can be retrieved by the method stated above, in real cases there are circumstances where convergence is very slow. For example, since the baseline must satisfy the boundary condition on both side, the trend of spectrum at both ends has a great impact to the shape of the iterating curve. In particular, for low or high pass signals of which the spectrum have a significant slope trend in low and high frequency regions, the iterating curve is prone to adhere the slope at both ends and continue to keep such inertia to some extent even when slipping out of those slope regions. This phenomenon brings us two bad effects in analysing signals: On one side, for low or high pass signals, their spectrum at both ends always exhibit half peaks since the scope of representable spectrum from FFT is always limited and the spectrum of those signals will inevitably hits the boundary, making the peak not complete and the estimation of center frequency inaccurate. On the other side, inertia effect will be harmful to the shape of the baseline, since in regions adjacent to low and high frequency bands, the function value of supporting baseline will have very large gap to the lower bound of the original spectrum due to the inertia of the downward trend. In order to circumvent such effect, we extend the original spectrum by forcing the spectrum to gently drop to the minimum of original spectrum and leave some extra flat intervals at both ends. This method not only take the advantage of trend effect, in that the curve will tend to keep horizontally at both ends which can make the peak be finally filtered out, but also complete the peaks at both ends without any harmful effects to the other regions as the experiment shows. Fig. 8 shows the evolution of spectrum baseline during the iteration with respect to the signal $y(t)=60t+60\cos 200\pi t$. This typical signal is a band-stop signal which shares both high and low pass feature. Note that if we don’t extend the spectrum, the baseline exhibits a sharp decreasing in the low frequency area and can not catch up the lower bound in the interval $[0.05,0.4]$ and the similar phenomenon also happens at the high-frequency area, even after 1500 iterations. This shows the necessity of proper extension to the original spectrum.
Figure 8: Iterations for finding the supporting baseline without extending the spectrum to the signal $y(t)=60t+60\cos 200\pi t$. The 500th(a), 1500th(b) iterations with no extension and the 500th(c), 1500th(d) iterations with extension are shown respectively. The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations.
4.3 Experiments on Central Frequencies
In order to prove that our algorithm can find frequencies close to the practical central frequencies for each IMF component, we compare our results from the above Example 1 to Example 5, with the results of the original VMD process with the same number of IMF, to keep away the impact to the performance of VMD from the initial frequency value that we provide to the following VMD procedure. We first get the frequency estimation from our algorithm and the result of pure VMD procedure with the same number of IMFs respectively, and then sort them in order, to observe the difference between the results of ours and the corresponding values from VMD.
Table 1: The evaluated center frequencies
Experiment Ours Original VMD Max Difference
1 0.0550 0.0496 0.0054
2 0.0300, 0.0575 0.0251, 0.0507 0.0068
3 0.0100, 0.0775 0.1550, 0.2075 0.0008, 0.0530 0.1510, 0.2 0.0245
4 0.0125, 0.0875 0.2475 6.46e-5, 0.0067 0.0803 0.1682
5 0.005,0.0775 0.1, 0.125 0.1550, 0.1800 0.2075, 0.2325 0.2575, 0.2825 0.3075 0.000845, 0.0747 0.0999, 0.1250 0.1510, 0.182 0.223, 0.25 0.2760, 0.303 0.339 0.0315
Table 1 shows the result, where in each experiment the evaluated center frequencies for both ours and the pure VMD are sorted, and the max difference is computed as $\max\limits_{i}|f_{i}-F_{i}|$, where $f$ stands for our estimated frequencies and $F$ pure VMD’s. As shown in Table 1, in most cases, our estimation to the center frequencies has low difference with respect to the pure VMD. It can be found that, in Experiment 4, the error between estimated center frequencies of ours and pure VMD’s are larger than the average, however, as analysed in Experiment 4, on one side, the original VMD, no matter given IMFs number 2 or 3, fails to achieve reconstruction error as small as ours, on the other side, as shown in Fig.S5.13a(c) even when both started at IMFs number 3, the original VMD seems produce replicated modes at the Gaussian-alike component, with almost the same amplitude and shape, while ours provide relatively independent components when given our initial estimation for the center frequencies, which seems to indicates that our estimation in this case is more reliable.
5 Conclusion
In this article, we proposed a globally convergent method that can automatically determine the number of intrinsic mode functions for Variational Mode Decomposition. For narrow-banded signals, this method does not need any prior knowledge about the scope of the IMFs number, and provide non-recursively schema to determine IMFs and corresponding central frequencies simultaneously. Experiments show that, our method can provide accurate number of IMFs and corresponding center frequencies even if the center frequencies are very close. Our implementation is performant in both accuracy and speed. Moreover, rigorous proof is given for the convergence of the method, making it theoretically reliable.
6 Acknowledgement
This study is supported by the National Natural Science Foundation of China under Grant No. 61773290 and the Fundamental Research Funds for the Central Universities (22120230311).
SUPPLEMENTARY MATERIAL
1 Introduction of VMD
The key idea of VMD is to decompose an input signal into a series number of sub-signals via variational optimization method in spectral domain[5]. It is assumed the frequency of each mode is almost compact around a center pulsation and the associated analytic signal is computed by means of Hilbert Transform to get the unilateral frequency spectrum. Then each mode’s frequency spectrum was shifted into baseband by multiplying a frequency modulation factor with respect to the estimated center frequency, transforming each band-pass mode into low pass, and the bandwidth can be estimated by $H^{1}$ Gaussian smoothness. Afterwards, they decompose the original signal into each modes by minimizing the sum of the bandwidth of each mode with the constraints that the summation of all modes is the original signal. In short, their key variation optimization can be listed as following:
$\min\limits_{u_{k},\omega_{k}}\left\{\sum\limits_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{\text{j}}{\pi t}\right)*u_{k}(t)\right]e^{-\text{j}\omega_{k}t}\right\|_{2}^{2}\right\}$
$\text{s.t.}\quad\sum\limits_{k}u_{k}=f$
so that each mode can be altered into an associated analytical signal by using Hilbert Transform and shifted into base-band for the following estimation of bandwidth using H1 Gaussian smoothness with the shifted demodulated components. To solve the problem, one can construct an augmented Lagrangian Function by introducing both quadratic penalty and Lagrangian Multipliers as follows:
$\mathcal{L}(u_{k},\omega_{k},\lambda)$ $=\alpha\sum_{k}\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}$
$\quad\quad\quad\quad\quad\quad\quad+\left\|f-\sum u_{k}\right\|_{2}^{2}+\langle\lambda,f-\sum u_{k}\rangle$
This problem can be solved by using ADMM(Alternate Direction Method of Multipliers), in which each component can be updated by solving the equivalent minimization problem:
$u_{k}^{n+1}=\underset{u_{k}\in\mathbb{R}}{\arg\min}\left\{\alpha\left\|\partial_{t}\left[\left(\delta(t)+\frac{j}{\pi t}\right)*u_{k}(t)\right]e^{-j\omega_{k}t}\right\|_{2}^{2}+\left\|f-\sum u_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
One can rewrite it in frequency domain form, so that convolution becomes multiplication for the convenience of the calculation in subsequent variational operations:
$\hat{u}_{k}^{n+1}=\underset{\hat{u}_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\alpha\left\|j\omega\left[1+\text{sgn}(\omega+\omega_{k})\right]\hat{u}_{k}(\omega+\omega_{k})\right\|_{2}^{2}+\left\|\hat{f}-\sum\hat{u}_{i}+\frac{\lambda}{2}\right\|_{2}^{2}\right\}.$
and by substitution of $\omega\rightarrow\omega+\omega_{k}$ to perform a translation and considering the Hermitian symmetry of spectrum with respect to the real signals, both terms can be transformed into half-space integrals with non-negative frequencies:
$\hat{u}_{k}^{n+1}=\underset{u_{k},\,\hat{u}_{k}=\hat{u}_{k}^{T}}{\arg\min}\left\{\int_{0}^{\infty}4\alpha(\omega-\omega_{k})^{2}|\hat{u}_{k}(\omega)|^{2}d\omega+2\int\left(\hat{f}-\sum\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)^{2}d\omega\right\}$
which can be evaluated by vanishing the first variation of positive frequencies for both signal components and center frequencies
$\hat{u}_{k}^{n+1}=\left(\hat{f}-\sum_{i\neq k}\hat{u}_{i}+\frac{\hat{\lambda}}{2}\right)\frac{1}{1+2\alpha(\omega-\omega_{k})^{2}}$
$\omega_{k}^{n+1}=\frac{\displaystyle\int_{0}^{\infty}\omega\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}{\displaystyle\int_{0}^{\infty}\left|\hat{u}_{k}(\omega)\right|^{2}d\omega}$
2 Finite Discrete Expression and Their Matrices
Here we show finite discrete expression and the corresponding the conversion matrices with respect to (11),(12). Since the finite difference expressions for $\alpha(x)$ and $g(x)$ are identical in form, the conversion matrices $\boldsymbol{G}^{(n)}$ and $\boldsymbol{A}^{(n)}$ also share the exactly the same structure
(S2.1) $g^{(1)}(x)\approx\cfrac{-g(x+2h)+8g(x+h)-8g(x-h)+g(x-2h)}{12h}$
(S2.2) $g^{(2)}(x)\approx\cfrac{-g(x+2h)+16g(x+h)-30g(x)+16g(x-h)-g(x-2h)}{12h^{2}}$
(S2.3) $g^{(3)}(x)\approx\cfrac{g(x+2h)-2g(x+h)+2g(x-h)-g(x-2h)}{2h^{3}}$
(S2.4) $g^{(4)}(x)\approx\cfrac{g(x+2h)-4g(x+h)+6g(x)-4g(x-h)+g(x-2h)}{h^{4}}$
(S2.5) $\boldsymbol{G}^{(1)}=\cfrac{1}{12h}\begin{bmatrix}1&-8&0&8&-1&0&\cdots&\cdots&\cdots&0\\ 0&1&-8&0&8&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-8&0&8&-1&0\\ 0&0&\cdots&\cdots&0&1&-8&0&8&-1\\ 12h&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&12h\\ -12&12&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-12&12\end{bmatrix}$
(S2.6) $\boldsymbol{G}^{(2)}=\cfrac{1}{12h^{2}}\begin{bmatrix}-1&16&-30&16&-1&0&\cdots&\cdots&\cdots&0\\ 0&-1&16&-30&16&-1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&16&-30&16&-1&0\\ 0&0&\cdots&\cdots&0&-1&16&-30&16&-1\\ 12h^{2}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&12h^{2}\\ -12h&12h&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-12h&12h\end{bmatrix}$
(S2.7) $\boldsymbol{G}^{(3)}=\cfrac{1}{2h^{3}}\begin{bmatrix}-1&2&0&-2&1&0&\cdots&\cdots&\cdots&0\\ 0&-1&2&0&-2&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&-1&2&0&-2&1&0\\ 0&0&\cdots&\cdots&0&-1&2&0&-2&1\\ 2h^{3}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&2h^{3}\\ -2h^{2}&2h^{2}&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-2h^{2}&2h^{2}\end{bmatrix}$
(S2.8) $\boldsymbol{G}^{(4)}=\cfrac{1}{h^{4}}\begin{bmatrix}1&-4&6&-4&1&0&\cdots&\cdots&\cdots&0\\ 0&1&-4&6&-4&1&0&\cdots&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&\cdots&\cdots&0&1&-4&6&-4&1&0\\ 0&0&\cdots&\cdots&0&1&-4&6&-4&1\\ h^{4}&0&0&\cdots&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&\cdots&0&0&h^{4}\\ -h^{3}&h^{3}&0&0&\cdots&\cdots&\cdots&0&0&0\\ 0&0&0&\cdots&\cdots&\cdots&0&0&-h^{3}&h^{3}\end{bmatrix}$
where the last four lines in matrices correspond to the boundary condition and the conversion matrices $\boldsymbol{A}^{(n)}$ have the same structure.
3 About Component-Wise Multiplication
Our component-wise multiplication is not necessarily the Hadamard production of matrices, in that we add some special rules in dealing with column vector, to say in detail, $\forall\boldsymbol{A}_{m\times n},\boldsymbol{v}_{m\times 1}$, we have
(S3.1) $\boldsymbol{v}\odot\boldsymbol{A}=\boldsymbol{A}\odot\boldsymbol{v}=\boldsymbol{A}*\boldsymbol{V}_{m\times n}$
where * stands for standard Hadamard production and each column of $\boldsymbol{V}$ is just a copy of $\boldsymbol{v}$. In this way, we can transform (13) into
(S3.2) $((2\boldsymbol{A^{(2)}\alpha})\odot\boldsymbol{G^{(2)}})g+((4\boldsymbol{A^{(1)}\alpha})\odot\boldsymbol{G^{(3)}})g+((2\boldsymbol{\alpha})\odot\boldsymbol{G^{(4)}})g=\boldsymbol{\beta-\lambda+\mu}$
By applying this rule we have $(\boldsymbol{A}\boldsymbol{a})\odot(\boldsymbol{B}\boldsymbol{b})=((\boldsymbol{A}\boldsymbol{a})\odot\boldsymbol{B})\boldsymbol{b}$ so that we can treat (13) as common matrix equation and get the solution.
4 Supplementary Material of the Proof
4.1 Proof of Convexity
Let $\mathcal{L}(g,\lambda,\mu,\alpha,\beta)=J_{1}(g(x))+J_{2}(g(x))$, where $J_{1}(g(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x$ and $J_{2}(g(x))=\displaystyle-\int_{\Omega}\beta(x)g(x)\text{d}x$. For $J_{1}(g(x))$, we have
(S4.1) $\lambda J_{1}(g_{1}(x))+(1-\lambda)J_{1}(g_{2}(x))-J_{1}[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]$
$=\lambda(1-\lambda)\int_{\Omega}\alpha(x)[g_{1}^{\prime\prime}(x)-g_{2}^{\prime\prime}(x)]^{2}\text{d}x\geq 0$
and $J_{2}(g(x))$ is obviously a linear funtional, so we have
(S4.2) $J[\lambda g_{1}(x)+(1-\lambda)g_{2}(x)]\leq\lambda J[g_{1}(x)]+(1-\lambda)J[g_{2}(x)]$
so that our objective functional is convex.
4.2 Proof of the Zero Duality Gap
Now we begin to prove that the duality gap is zero. Let
(S4.3) $\mathcal{O}(g(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x-\int_{\Omega}\beta(x)g(x)\text{d}x$
be the objective functional, and construct Lagrangian functional
$\mathcal{L}(g(x),\lambda(x),\mu(x))=\displaystyle\int_{\Omega}\alpha(x)g^{\prime\prime 2}(x)\text{d}x$ $-\int_{\Omega}\beta(x)g(x)\text{d}x$
(S4.4) $+\int_{\Omega}\lambda(x)[g(x)-f(x)]\text{d}x-\int_{\Omega}\mu(x)g(x)\text{d}x$
define dual functional as
(S4.5) $\mathcal{G}(\lambda(x),\mu(x))=\inf\limits_{g(x)\in\mathcal{D}}\mathcal{L}(g(x),\lambda(x),\mu(x))$
then on one side, we must have
(S4.6) $\mathcal{G}(\lambda^{*}(x),\mu^{*}(x)=\inf\limits_{g(x)}\mathcal{L}(g(x),\lambda^{*}(x),\mu^{*}(x))\leq\inf\limits_{g(x)}\mathcal{O}(g(x))=\mathcal{O}(g^{*}(x))$
where * stands for the optimum. On the other hand, one can construct general sets
(S4.7) $\mathcal{S}=\{(p(x),q(x),o)|g(x)-f(x)\leq p(x),-g(x)\leq q(x),\mathcal{O}(g(x))\leq o\}$
which forms an epi-graph to (2). thus
(S4.8) $\forall\xi(x)=\lambda(p_{1}(x),q_{1}(x),o_{1})+(1-\lambda)(p_{2}(x),q_{2}(x),o_{2})$
one can always find
(S4.9) $\eta(x)=\lambda g_{1}(x)+(1-\lambda)g_{2}(x)$
that makes
[…]
hold and thus $\mathcal{S}$ is a convex set. It is obvious that $(0,0,\mathcal{O}(g^{*}(x)))\in\partial S$, so according to the Support Theorem of Convex Sets there must be a general hyperplane that on one side pass through $(0,0,\mathcal{O}(g^{*}(x)))\in\partial S$ and on the other side ensures any point in $\mathcal{S}$ locates above the hyperplane. That is, there exists $\lambda(x),\mu(x)$ that makes
(S4.11) $\int_{\Omega}\lambda(x)p(x)\text{d}x+\int_{\Omega}\mu(x)q(x)\text{d}x\geq\mathcal{O}(g^{*}(x))-o$
hold for all $(p(x),q(x),o)$ in $\mathcal{S}$. Note that at least $\forall p(x),q(x)\geq 0$, $\mathcal{O}(g^{*}(x))\geq o$, so that the right end in this case is always non-negative. To summarize all above, we must have $\lambda(x)\geq 0$ and $\mu(x)\geq 0$, which indicates we can find a solution in dual space making the dual-gap vanish.
Next we prove the convexity of our feasible domain. The feasible set is $0\leq g(x)\leq f(x)$, so that $\forall g_{1}(x),g_{2}(x)\in[0,f(x)]$, we have $\lambda g_{1}(x)+(1-\lambda)g_{2}(x)\in[0,f(x)],\forall\lambda\in[0,1]$, which means the feasible set is also convex.
5 Experiment Detail
In this section we begin to provide some details of our experiments as complementary.
Figure S5.1: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$. The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iteration are respectively shown.
Figure S5.2: The gap-elimination course for the found spectrum supporting line with series $y(t)=100\sin 20\pi t$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
Figure S5.3: The decomposition result of $y(t)=100\sin 20\pi t$ from the subsequent VMD, using the found number of modes and center frequency as initialization. The subfigure from top to bottom are respective the source signal, the decomposed mode(only one) and the residual.
Figure S5.4: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$. The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iterations are respectively shown.
Figure S5.5: The gap-elimination course for the found spectrum supporting line with series $y(t)=10\cos 10\pi t+20\sin 20\pi t$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
Figure S5.6: The decomposition result of $y(t)=10\cos 10\pi t+20\sin 20\pi t$ from the subsequent VMD, using the found number of modes and center frequency as initialization. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are two) and the residual.
Figure S5.7: The iteration for finding the supporting baseline of FFT with respect to time series (45). The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown. Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out.
Figure S5.8: The gap-elimination course for the found spectrum supporting line with series (45). (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom.
Figure S5.9: The decomposition result of (45) from the subsequent VMD, using the found number of modes and center frequency as initialization. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are four) and the residual.
Figure S5.10: The iterations for finding the supporting baseline of FFT with respect to time series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ are shown. The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown.
Figure S5.11: The noise-elimination course for the found spectrum supporting line with series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
Figure S5.12: (a) and (b)
Figure S5.13: The decomposition result (a) of $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ from the subsequent VMD, using the found number of modes and center frequency as initialization, compared with original VMD results with IMFS number 2 (b) and 3 (c) respectively. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual. As compared with (b) and (c), for original VMD, both the results has reconstruction error twice as large as ours.
Figure S5.14: The iteration for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The 500th(a), 1000th(b), 1500th(c) iterations are respectively shown.
Figure S5.15: The gap-elimination course for the found spectrum supporting line with series $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
Figure S5.16: The decomposition result of $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$ from the subsequent VMD, using the found number of modes and center frequency as initialization. The sub-figure from top to bottom are respectively the source signal, the decomposed modes(there are three) and the residual.